# ExLM: Rethinking the Impact of [MASK] Tokens in Masked Language Models

### v2

@@ 4.3 Pre-training of ExLM: States Alignment
RoBERTa 84.7/84.9 91.3 91.4 93.5 61.3 75.5 88.7 89.5 85.2 78.3 81.5
ExLM 87.5/87.3 92.2 93.4 93.9 64.6 77.2 89.6 90.5 86.2 81.3 84.6
ExLM-Large 88.7/88.6 92.7 93.8 94.5 65.3 78.1 90.6 91.2 87.0 82.9 86.1
5 Experimental Results
We evaluate the performance of the ExLM model in both text modeling and SMILES modeling tasks. SMILES is a textual representation of molecular information, and a more detailed explanation can be found in Appendix D.
We pre-train ExLM on data of text and SMILES. Then we fine-tune and evaluate ExLM across diverse benchmarks and verify the contribution of each component through ablation experiments. Finally, a visualization analysis is included to explain the advantages of unified modeling.
5.1 Results on Molecular Representation Learning
Pre-training.
Fine-tuning.
Results.
5.2 Results on Textual Representation Learning
Pre-training.
For textual pre-training, we adopt the English Wikipedia and BookCorpus datasets (Devlin, 2018) as the pre-training dataset. The model size of ExLM is consistent with BERT base (Devlin, 2018). We also train ExLM-Large with the same size as BERT large. For more details about pre-training settings, please see Appendix N.
Fine-tuning.
Results.
As shown in Table 2, we evaluate ExLM on the GLUE and SQuAD 2.0 development sets. ExLM achieves the best performance on SQuAD 2.0 and 7 GLUE tasks, and also performs very closely to the best baseline model on the remaining one GLUE task. Furthermore, ExLM demonstrates significantly higher average performance compared to other baseline models. Notably, RoBERTa in the table uses the same pre-training data and settings as ExLM, yet ExLM shows significant improvements. This demonstrates the superior performance and effectiveness of ExLM. ExLM-Large also achieves obvious performance improvements compared to ExLM due to the increased scale.
@@ Results.
ExLM w/o 2D RoPE 84.6 91.1 91.6 56.7 81.0
ExLM w/o Transitions 83.8 90.9 91.1 55.6 80.4
Vanila MLM 83.6 90.0 90.3 54.7 79.6
Vanila MLM++ 84.4 90.6 91.5 56.2 80.7
ExLM 85.1 91.8 92.2 57.5 81.6
5.3 Analytical Experiments of ExLM
We validate ExLM’s effectiveness through ablation studies and explore the impact of expanded states on performance. A case study further demonstrates its good ability to capture rich semantic information. Additional analyses on training efficiency and mask ratio are also provided in Appendix Q.
Table 4: Impact of the number of expanded states $k$. Both overly large and small values of $k$ negatively affect ExLM’s performance, and $k=4$ is a relatively optimal choice.
[…]
Vanila MLM 83.6 90.0 90.3 54.7 79.6
ExLM-$k=1$ 83.5 90.1 90.3 54.8 79.6
ExLM-$k=2$ 84.3 91.3 91.6 57.0 81.0
ExLM-$k=4$ 85.1 91.8 92.2 57.5 81.6
ExLM-$k=8$ 84.1 90.7 91.3 56.9 80.8
@@ 5.3 Analytical Experiments of ExLM
The impact of the number of expanded states.
We explore the impact of the number of expanded states, $k$, on ExLM’s performance, with results shown in Table 4. When $k=1$, ExLM degenerates into an MLM, yielding performance identical to that of MLM, as expected. As $k$ increases, performance improves due to enhanced semantic modeling ability. However, when $k$ becomes too large (e.g., $k=8$), performance slightly declines due to the excessive expanded states of [MASK] tokens in the input context and redundant nodes in the DAG. Thus, with a mask ratio of $15\%$, $k=4$ proves to be the optimal choice.
Figure 7: Case study. We visualize the model’s predictions when the input is ”This is [MASK], and I’m very [MASK] to see this.” ($k=4$). The yellow nodes correspond to the expanded states of the first [MASK] token, while the brown nodes correspond to those of the second [MASK] token. The weights in the figure represent the transition probability between two nodes in the DAG, and the axis labels show the top-1 predicted word for each node.
@@ 6 Conclusions
In this paper, we analyze the impact of the semantics corruption caused by [MASK] tokens on MLMs, showing that it has a more significant impact than the unreal token problem, offering a new perspective on understanding MLMs. Based on our analysis, we propose ExLM, which enhances the model’s semantic modeling ability through two key designs: States Expansion and Dependency Capture. These designs reduce the impact of multimodal issues and increase semantic richness. We demonstrate ExLM’s strong performance in both text and SMILES modeling scenarios. Ablation studies and case study further validate its effectiveness.
@@ Acknowledgements
We would like to thank all other members from Dlib in PKU for their support and insightful discussions on the project. This paper is partially supported by National Natural Science Foundation of China with Grant Number 62276002.
Impact Statement
This paper is the first work to systematically study the impact of [MASK] tokens on MLM from the perspectives of unreal tokens and semantics corruption, and proposes effective solutions to these issues. These analytical results provide valuable insights for designing better language representation learning models. Furthermore, MLM-based sequence representation learning models have already achieved significant success in various domains, including text, images, video, and scientific data. These models have also shown promising results in applications such as recommendation systems and protein design. Therefore, the methods proposed in this paper have the potential to drive further advancements in these fields, and the approach also holds significant application potential. However, we also acknowledge that this work inherits the potential negative impacts of existing MLMs, such as the possibility of being used to generate false information or design and manufacture molecules with biological hazards.
@@ Appendix A Related Works
Currently, language models serve as an important method in the field of self-supervised representation learning and achieve excellent performance across a wide range of tasks. The ELMo model (Peters et al., 2018) first introduces a self-supervised language representation learning model based on bidirectional LSTM (Hochreiter, 1997). Later, the pre-trained masked language model based on the transformer architecture, BERT (Devlin, 2018), is proposed in the field of text modeling. BERT uses a bidirectional transformer encoder to extract meaningful semantic representations from large amounts of text data and applies them to various downstream tasks such as text understanding (Devlin, 2018; Liu, 2019) and text correction (Zhang et al., 2020; Dai et al., 2022). With the help of the transformer architecture (Vaswani et al., 2017) and MLM training approach, this model achieves great success in the text domain. Following this, a large number of follow-up works aim to improve these bidirectional encoder-based masked language models (Liu, 2019; Lan, 2019; Ke et al., 2020; Clark, 2020; Joshi et al., 2020; He et al., 2020; Shin et al., 2020; Du et al., 2021; Fu et al., 2022; Wettig et al., 2022; Meng et al., 2022, 2023; He et al., 2022b). Moreover, MLM also succeeds in various sequence representation learning tasks, including images (Bao et al., 2021; Xie et al., 2022; He et al., 2022a), videos (Tong et al., 2022; Wang et al., 2022), codes (Feng et al., 2020; Guo et al., 2020; He et al., 2021), small molecules(Wang et al., 2019; Chithrananda et al., 2020; Pan, 2023; Feng et al., 2024; Yang et al., 2024; Zheng et al., 2024a), proteins (Elnaggar et al., 2021; Lin et al., 2022, 2023; Zheng et al., 2024b; Hayes et al., 2025), single-cell data (Yang et al., 2022; Theodoris et al., 2023), and others.
@@
In addition to language models trained with the MLM pretraining approach, there are other types of language models used for representation learning tasks, such as autoregressive language models (Wang et al., 2023; Jiang et al., 2023; Springer et al., 2024; Ren et al., 2024; Li & Li, 2024), diffusion language models (Wang et al., 2024a, b), and unified language models (Dong et al., 2019; Bao et al., 2020; Du et al., 2021). These models have demonstrated their effectiveness in representation learning across various tasks. However, overall, MLM is one of the most commonly used approaches for constructing language models for representation learning tasks and is a typical representative of this type of model. Given the widespread impact and application of MLM, studying the corrupted semantics problem in MLM is of great significance, and it may also play a positive role in advancing the future development of MLM in different fields.
@@
Due to the widespread use of MLM and the important role of the [MASK] tokens in the MLM training process, several previous works have studied the impact of the [MASK] tokens on MLM. Clark (2020) proposed the ELECTRA model, which is an improvement on MLM. In this model, the [MASK] token is not introduced as part of the model’s context during pretraining. This work also suggests that this training approach can address the inconsistency issue between pretraining and downstream task fine-tuning caused by the [MASK] unreal tokens. Liao et al. (2022) further studied the impact of the [MASK] unreal tokens on MLM models and pointed out that removing [MASK] tokens from the model’s shallow representations during training does not affect MLM performance. Meanwhile, Meng et al. (2023) explored the representation deficiency problem caused by [MASK] unreal tokens from a theoretical perspective and proposed a MAE-based language model training approach. This approach improves the model’s representation learning ability by removing the [MASK] tokens from the model input to avoid the unreal tokens issue. Additionally, Wettig et al. (2022) and Liao et al. (2022) also explored the impact of higher mask ratio on MLM performance. While these works have studied the effects of the [MASK] tokens, they focus only on the unreal tokens aspect and lack an investigation into the model’s behavior and performance from the perspective of semantic corruption. This work aims to fill this gap by analyzing the problem from the perspective of both the semantic corruption issue and the unreal tokens issue.
@@ B.1 Pre-training Configuration
To ensure the comparability of the results, all other pre-training hyperparameters in the MLM across different experimental groups are exactly the same, except for the Max sequence length parameter. For more pre-training hyper-parameters, please refer to Table 5.
@@ Appendix H Dynamic Programming for Efficient Alignment
Naively summing over all paths $\mathbf{A}\in\Gamma$ can be prohibitively expensive, since the number of possible paths grows exponentially. To solve this, we use a dynamic programming (DP) approach, which is inspired by methods for connectionist temporal classification (Graves et al., 2006) and DA-Transformer (Huang et al., 2022).
@@
By repeatedly computing $f_{i,u}$ for all $i$ and $u$, we ultimately obtain $f_{M,L}$ (where $M$ is the length of the target $\mathbf{Y}$, and $L$ indexes the final state). The final training objective is:
@@ I.2 Gradient Calculations
Gradient computation is essential for optimizing model parameters during training. The algorithm calculates gradients with respect to both emission probabilities $\mathbf{P}_{u}(y_{i})$ and transition scores $\mathbf{E}_{v,u}$. In the following formulas, $M$ denotes the final position in the target sequence (i.e., the last index along the target length dimension), while grad_output refers to the upstream gradient propagated from the overall loss function.
[…]
where $\log(\sum_{u^{\prime}}\alpha_{M,u^{\prime}})$ represents the log-likelihood of the entire sequence, computed by summing the forward probabilities $\alpha_{M,u^{\prime}}$ over all possible final states $u^{\prime}$ at the final position $M$, grad_output is the gradient passed down from the final training objective, indicating how the overall loss changes with respect to the intermediate gradients.
@@
[…]
This gradient aggregates the impact of each transition from state $v$ to state $u$ across all relevant positions $i$. Similar to the emission gradient, it subtracts $\log(\sum_{u^{\prime}}\alpha_{M,u^{\prime}})$ to ensure proper normalization based on the total sequence likelihood, and it is scaled by grad_output.
@@ I.3 Time Complexity Analysis
where bsz is batch size, tarlen is target sequence length, prelen is the number of states, translen is the maximum number of transitions per state.
@@ I.4 Space Complexity Analysis
where bsz is batch size, tarlen is target sequence length, prelen is the number of states, translen is the maximum number of transitions per state.
@@
We evaluate our approach against various supervised learning and pre-training baselines, including both SMILES-based and 3D molecular pre-training models. The supervised methods include D-MPNN (Yang et al., 2019) and AttentiveFP (Xiong et al., 2019), both of which are based on graph neural networks (GNNs). For 2D and 3D molecular pre-training, we consider baseline methods: N-gram (Liu et al., 2019), GROVER (Rong et al., 2020), GraphMVP (Liu et al., 2021), Mole-BERT (Xia et al., 2023), and 3D InfoMax (Stärk et al., 2022). For a fair comparison, we train a SMILES model based on MLM pre-training, referred to as SMILES-BERT, using the same training data, model architecture, and training hyperparameters as ExLM.
@@
ESOL 1 Regression 902/113/113 Water solubility
FreeSolv 1 Regression 513/64/64 Hydrogen free energy
Lipo 1 Regression 3,360/420/420 Octanol/water distribution ratio, coefficient
@@
We use Spearman correlation for STS, Matthews correlation for CoLA, and accuracy for the other tasks as the metrics on GLUE.
@@
We implement ExLM in two configurations: a base model and a large model. The base ExLM consists of 12 stacked Transformer layers, each with 12 attention heads. The model dimension and feedforward dimension of each Transformer layer are $768$ and $3{,}072$, respectively, resulting in a total of $128$M parameters. The large ExLM model uses $24$ Transformer layers with $16$ attention heads per layer. The model dimension and feedforward dimension are increased to $1,024$ and $4,096$, respectively, with a total parameter count of $354$M. We use Adam (Kingma & Ba, 2014) optimizer and polynomial learning rate scheduler to train ExLM, and we set the learning rate as 5e-4 and warmup steps as $10$K. The total training steps are $125$K and each batch has $2048$ samples at maximum. We also implement the ExLM model using the Fairseq library. For more pre-training hyper-parameters, please refer to Table 10.
@@
Figure 10: We visualize the model’s predictions when the input is ”This is [MASK], and I’m very [MASK] to see this.” ($k=4$). The yellow nodes represent the expanded states corresponding to the first [MASK] token, while the brown nodes represent the expanded states corresponding to the second [MASK] token. The x-axis and y-axis show the top-3 predicted word for each graph node.
@@
  • ExLM successfully represents the different choices of each [MASK] token using distinct expanded states. Since $k=4$, the model predicts four different choices for both [MASK] tokens in the input. These four choices typically represent different aspects of the semantics. For example, the first [MASK] token’s choices include three categories: a noun (e.g., it, news), negative emotion words (e.g., terrible, awful), and positive emotion words (e.g., amazing, beautiful, wonderful), each represented by a different expanded state. This effectively avoids semantic ambiguity in the context, i.e., the multimodal issue. Similarly, the second [MASK] token’s choices include positive emotions (e.g., glad, happy) and negative emotions (e.g., sorry), and these different types of semantic information are also represented by distinct expanded states.
@@
  • ExLM typically has lower uncertainty for [MASK] tokens that appear later. As shown in 10, the word probability distributions of the four expanded states corresponding to the first [MASK] token are often more evenly distributed, meaning that the differences of word probabilities across different words in each state are not very large. However, for the four expanded states corresponding to the second [MASK] token, the model usually predicts one word with a probability significantly higher than the others. This phenomenon occurs due to the directionality of the DAG. Specifically, in ExLM’s DAG, transitions only occur from earlier states to later states, and the choices of earlier states have a greater impact on the final result, which means that the earlier states tend to have higher uncertainty.
@@
Q.1 The Impact of Mask Ratio $p$ on ExLM
We train ExLM with different mask ratios $p$ and numbers of expanded states $k$ using the hyperparameters from Table 5, and compare its performance with that of MLM. The results are shown in Table 12. From these results, we observe that although an increase in mask ratio significantly degrades MLM performance, ExLM maintains strong performance even under higher mask ratios. This demonstrates that despite the more severe multimodal issues with a higher mask ratio, ExLM is better at handling multimodal issues in the input context compared to MLM, and it can still perform well under high mask ratios. In this case, the performance advantage of ExLM over MLM is further amplified. Furthermore, we find that ExLM requires more expanded states to effectively model the semantically multimodal and ambiguous contextual information resulting from the higher mask ratio. Therefore, appropriately increasing $k$ at higher mask ratios can still help the model improve its performance. However, an excessively high mask ratio and too many expanded states (higer $k$) can result in an excessive number of expanded states of [MASK] tokens in the input context, limiting further performance gains.
Table 12: Performance comparison of MLM and ExLM with different mask ratios $p$ and numbers of expanded states $k$. ExLM demonstrates stronger modeling capability for contexts with a high mask ratios compared with MLM, and increasing $k$ further improves the model’s performance under such high mask ratios inputs.
[…]
$p=15\%$ Vanila MLM 83.6 90.0 90.3 54.7 79.6
$p=15\%$ ExLM-$k=2$ 84.3 91.3 91.6 57.0 81.0
ExLM-$k=4$ 85.1 91.8 92.2 57.5 81.6
ExLM-$k=8$ 84.1 90.7 91.3 56.9 80.8
$p=38.7\%$ Vanila MLM 83.3 90.0 90.2 53.3 79.2
$p=38.7\%$ ExLM-$k=2$ 84.2 91.4 91.6 56.5 80.9
ExLM-$k=4$ 84.3 91.6 92.1 57.0 81.3
ExLM-$k=8$ 84.0 91.2 91.1 56.2 80.6
We also conduct an entropy analysis to compare the severity of the multimodal problem under different mask ratios $p$ in the input context for ExLM and MLM. The results are shown in Figure 12. The results reveal that ExLM has significantly lower entropy in its prediction distribution than MLM, indicating lower uncertainty in dealing with ambiguous contexts. This further proves that ExLM significantly mitigates the negative impact of multimodal issues. Additionally, as $k$ increases, the uncertainty of ExLM decreases, suggesting that increasing $k$ expands the model’s semantic space, helping to alleviate the negative effects of multimodal issues. Moreover, under higher mask ratios, ExLM shows a more noticeable reduction in uncertainty compared to MLM, explaining why ExLM still maintains relatively strong performance under high mask ratios.
@@
We compare the training efficiency differences between MLM and ExLM ($k=4$) with the same pre-training configuration. Both models are trained on two Tesla A100 80G GPUs under the hyperparameters from Table 5, and their training cost statistics are shown in Table 13. From this table, we can know that the training cost of ExLM ($k=4$) is approximately 2.2 times that of the MLM with the same pre-training configuration. To compare the performance differences between MLM and ExLM ($k=4$) under equal training costs, we increase the training steps of MLM from $50{,}000$ to $110{,}000$, and refer to this model as Vanilla MLM++. The performance of Vanilla MLM++ is shown in Table 3, where we can see that although increasing the training steps significantly improves MLM performance, it still remains lower than that of ExLM. This further demonstrates that ExLM achieves stronger performance under the same training cost.

### v3

@@ Results.
RoBERTa 85.9/85.8 91.6 92.3 93.7 64.3 75.5 88.7 89.5 85.2 78.3 81.5
ExLM 86.9/86.7 92.0 93.1 93.9 64.6 78.8 89.6 90.5 86.2 82.0 84.6
ExLMLARGE 87.8/87.5 92.2 93.8 94.5 65.3 79.1 90.4 91.2 86.9 82.6 85.0
5.3 Analytical Experiments of ExLM
We validate ExLM’s effectiveness through ablation studies and explore the impact of expanded states and mask ratios on its performance. A case study also demonstrates its good ability to capture rich semantic information. Efficiency and entropy analysis of ExLM are provided in Appendix Q.
Table 4: Performance comparison of MLM and ExLM with different mask ratios $p$ and numbers of expanded states $k$. ExLM demonstrates stronger modeling capability for contexts with a high mask ratios compared with MLM, and increasing $k$ further improves its performance under high mask ratios inputs.
@@ 5.3 Analytical Experiments of ExLM
$15\%$ Vanilla MLM 83.6 90.0 90.3 54.7 79.6
$15\%$ ExLM-k=$2$ 84.6 91.3 91.1 56.7 80.9
ExLM-k=$4$ 85.1 91.4 91.3 57.6 81.4
ExLM-k=$8$ 84.4 91.0 90.9 56.9 80.8
$38.7\%$ Vanilla MLM 83.3 90.0 90.2 53.3 79.2
$38.7\%$ ExLM-k=$2$ 84.1 91.4 91.0 55.4 80.5
ExLM-k=$4$ 84.9 91.6 91.2 57.0 81.2
ExLM-k=$8$ 84.3 90.9 90.7 56.8 80.7
@@
The impact of $k$ and $p$.
We explore the impact of the number of expanded states $k$ and mask ratios $p$ on ExLM’s performance, with results shown in Table 4. As $k$ increases, performance improves due to enhanced semantic modeling ability. However, when $k$ becomes too large (e.g., $k=8$), performance slightly declines due to the excessive expanded states of [MASK] tokens in the input context and redundant nodes in the DAG. Thus, with a mask ratio of $15\%$, $k=4$ proves to be the optimal choice. We further verify the impact of the mask ratio $p$ on the performance of ExLM, with the results shown in Table 4. These results show that a higher mask ratio $p$ has less impact on ExLM’s performance compared to MLM, highlighting the enhanced semantic modeling capability of ExLM. Furthermore, ExLM needs more expanded states to model the multimodal and ambiguous context induced by a higher mask ratio. Thus, increasing $k$ appropriately at higher mask ratios enhances its performance. However, an excessively high mask ratio and too many expanded states (a higher $k$) can result in a large number of expanded states of [MASK] tokens in the input context, limiting further performance gains of ExLM.
Figure 7: Case study. We visualize the model’s predictions when the input is “This is [MASK], and I’m very [MASK] to see this.” ($k=4$). The yellow nodes correspond to the expanded states of the first [MASK] token, while the brown nodes correspond to those of the second [MASK] token. The weights in the figure represent the transition probability between two nodes in the DAG, and the axis labels show the top-1 predicted word for each node.
@@ 6 Conclusions
In this paper, we analyze the impact of the semantics corruption caused by [MASK] tokens on MLMs, showing that it has a more significant impact than the unreal token problem, offering a new perspective on understanding MLMs. Based on our analysis, we propose ExLM, which enhances the model’s semantic modeling ability through two key designs: States Expansion and Dependency Capture. These designs reduce the negative impact of the semantic multimodality on the model. We demonstrate ExLM’s strong performance in both text and SMILES modeling scenarios. Ablation studies and case study also validate its effectiveness and efficiency.
Impact Statement
This paper is the first work to systematically study the impact of [MASK] tokens on MLM from the perspectives of unreal tokens and semantics corruption, and proposes effective solutions to these problems. These analytical results provide valuable insights for designing better language representation learning models. Furthermore, MLM-based sequence representation learning models have already achieved significant success in various domains, including text, images, video, and scientific data. These models have also shown promising results in applications such as recommendation systems and protein design. Therefore, the methods proposed in this paper have the potential to drive further advancements in these fields, and the approach also holds significant application potential. However, we also acknowledge that this work inherits the potential negative impacts of existing MLMs, such as the possibility of being used to generate false information or design and manufacture molecules with biological hazards.
@@ Acknowledgements
We would like to thank Jie Zhu from University of Oxford and Liang Ding from the JD Explore Academy for their insightful discussions on the project. We also thank all other members from Dlib in PKU for their valuable feedback given during the internal discussions. This paper is partially supported by National Natural Science Foundation of China with Grant Number 62276002.
@@ Appendix A Related Works
Currently, language models serve as an important role in the self-supervised representation learning field and achieve excellent performance across a wide range of tasks. The ELMo model (Peters et al., 2018) first introduces a self-supervised language representation learning model based on bidirectional LSTM (Hochreiter, 1997). Later, the pre-trained masked language model based on the transformer architecture, BERT (Devlin, 2018), is proposed in the field of text modeling. BERT uses a bidirectional transformer encoder to extract meaningful semantic representations from large amounts of text data and applies them to various downstream tasks such as text understanding (Devlin, 2018) and text correction (Zhang et al., 2020; Dai et al., 2022). With the help of the transformer architecture (Vaswani et al., 2017) and the MLM training approach, this model achieves great success in the text domain. Following this, a large number of follow-up works aim to improve these bidirectional encoder-based masked language models (Liu, 2019; Lan, 2019; Ke et al., 2020; Joshi et al., 2020; He et al., 2020; Shin et al., 2020; Du et al., 2021; Fu et al., 2022; Wettig et al., 2022; Meng et al., 2022, 2023; He et al., 2022b). Moreover, MLM also succeeds in various representation learning tasks, including images (Bao et al., 2021; Xie et al., 2022; He et al., 2022a), videos (Tong et al., 2022; Wang et al., 2022), codes (Feng et al., 2020; Guo et al., 2020; He et al., 2021), small molecules(Wang et al., 2019; Chithrananda et al., 2020; Ross et al., 2022; Pan, 2023; Feng et al., 2024; Yang et al., 2024; Zheng et al., 2024a), proteins (Elnaggar et al., 2021; Lin et al., 2022, 2023; Zheng et al., 2024b; Hayes et al., 2025), and single-cell data (Yang et al., 2022; Theodoris et al., 2023).
@@
In addition to language models trained with the MLM pre-training approach, there are other types of language models used for representation learning tasks, such as autoregressive language models (Wang et al., 2023; Jiang et al., 2023; Springer et al., 2024; Ren et al., 2024; Li & Li, 2024), diffusion language models (Wang et al., 2024a, b), and unified language models (Dong et al., 2019; Bao et al., 2020; Du et al., 2021). These models have demonstrated their effectiveness in representation learning across various tasks. However, overall, MLM is one of the most commonly used approaches for constructing language models for representation learning tasks and is a typical representative of this type of model. Zhong et al. (2023a) also points out that the MLM performs better than LLM in handling paraphrase and similarity tasks. Given the widespread impact and application of MLM, studying the corrupted semantics problem in MLM is of great significance, and it may also play a positive role in advancing the future development of MLM in different fields.
@@
Due to the widespread use of MLM and the important role of the [MASK] tokens in the MLM training process, several previous works have studied the impact of the [MASK] tokens on MLM. Clark (2020) proposed the ELECTRA model, an improvement on MLM, where the [MASK] token is not included in the context during pre-training. This approach helps address the inconsistency between pre-training and downstream task fine-tuning caused by [MASK] unreal tokens. Liao et al. (2022) further studied the impact of the [MASK] unreal tokens on MLM models and pointed out that removing [MASK] tokens from the model’s shallow representations during training does not affect MLM performance. Meanwhile, Meng et al. (2023) explored the representation deficiency problem caused by [MASK] unreal tokens from a theoretical perspective and proposed a MAE-based language model training approach. This approach improves the model’s representation learning ability by removing the [MASK] tokens from the model input to avoid the unreal tokens problem. Additionally, Wettig et al. (2022) and Liao et al. (2022) also explored the impact of higher mask ratio on MLM performance. While these works have studied the effects of the [MASK] tokens, they focus only on the unreal tokens aspect and lack an investigation into the model’s behavior from the perspective of semantic corruption. In contrast, Zhong et al. (2023b) points out that dropping [MASK] tokens causes semantic loss, affecting performance on semantic-intense tasks like RTE (Dagan et al., 2005). This finding further highlights the negative impact of token dropping on the model’s semantic modeling capability, emphasizing the importance of addressing the semantic loss problem. However, this work only analyzes the additional semantic loss caused by the token dropping process compared to the token masking process, and does not analyze the semantic loss caused by the token masking process itself. Therefore, its conclusion can not generalize to MLMs. This work aims to fill this gap by analyzing MLM’s behavior from the perspective of both the semantic corruption problem and the unreal tokens problem.
@@ B.1 Pre-training Configuration
To ensure the comparability of the results, all other pre-training hyperparameters in the MLM across different experimental groups are exactly the same, except for the Max sequence length parameter. In addition to the Repeated MLM experiment, we also used the same pre-training parameters in the ablation experiments to reduce pre-training costs. For more pre-training hyper-parameters, please refer to Table 5.
@@ Appendix H Dynamic Programming for Efficient Alignment
Naively summing over all paths $\mathbf{A}\in\Gamma$ can be prohibitively expensive, since the number of possible paths grows exponentially. To solve this, we use a dynamic programming (DP) approach, which is also adopted by the training algorithms for connectionist temporal classification (Graves et al., 2006; Saharia et al., 2020) and DA-Transformer (Huang et al., 2022; Shao et al., 2022; Huang et al., 2023, 2024).
@@
By repeatedly computing $f_{i,u}$ for all $i$ and $u$, we ultimately obtain $f_{M,L}$ (where $M$ is the length of the target sequence $\mathbf{Y}$ (equal to the number of masked tokens), and $L$ indexes the final state.). The final training objective is:
@@ I.2 Gradient Calculations
Gradient computation is essential for optimizing model parameters during training. The algorithm calculates gradients with respect to both emission probabilities $\mathbf{P}_{u}(y_{i})$ and transition scores $\mathbf{E}_{v,u}$. In the following formulas, $M$ denotes the final position in the target sequence (i.e., the last index along the target length dimension).
$\nabla\mathbf{P}_{u}(y_{i})=\frac{\alpha_{i,u}\times\beta_{i,u}}{\sum_{u}\alpha_{M,u}}$
where $\sum_{u}\alpha_{M,u}$ represents the likelihood of the entire sequence, computed by summing the forward probabilities $\alpha_{M,u^{\prime}}$ over all possible final states $u^{\prime}$ at the final position $M$. This gradient indicates how changes in the emission probability for state $u$ at position $i$ affect the log-likelihood of the sequence.
@@
$\nabla\mathbf{E}_{v,u}=\frac{\alpha_{i-1,v}\times\beta_{i,u}}{\sum_{u}\alpha_{M,u}}$
This gradient measures the effect of changes in the transition score between states $v$ and $u$ on the overall log-likelihood of the sequence.
@@ I.3 Time Complexity Analysis
where bsz is batch size, tarlen is target sequence length, prelen is the number of states, translen is the maximum number of transitions per state (usually equals to prelen).
@@ I.4 Space Complexity Analysis
where bsz is batch size, tarlen is target sequence length, prelen is the number of states, translen is the maximum number of transitions per state (usually equals to prelen).
@@
We evaluate our approach against various supervised learning and pre-training baselines, including both SMILES-based and 3D molecular pre-trained models. The supervised methods include D-MPNN (Yang et al., 2019) and AttentiveFP (Xiong et al., 2019), both of which are based on graph neural networks (GNNs). For 2D and 3D molecular pre-training, we consider baseline methods: N-gram (Liu et al., 2019), GROVER (Rong et al., 2020), GraphMVP (Liu et al., 2021), Mole-BERT (Xia et al., 2023), and 3D InfoMax (Stärk et al., 2022). For a fair comparison, we train a SMILES model based on MLM pre-training, referred to as SMILES-BERT, using the same training data, model architecture, and training hyperparameters as ExLM.
@@
We use Spearman correlation for STS, Matthews correlation for CoLA, and accuracy for MNLI, QNLI, RTE and SST-2 as the metrics on GLUE.
@@
We implement ExLM in two configurations: a base model and a large model. The base ExLM consists of 12 stacked Transformer layers, each with 12 attention heads. The model dimension and feedforward dimension of each Transformer layer are $768$ and $3{,}072$, respectively, resulting in a total of $128$M parameters. The large ExLM model uses $24$ Transformer layers with $16$ attention heads per layer. The model dimension and feedforward dimension are increased to $1,024$ and $4,096$, respectively, with a total parameter count of $361$M. We use Adam (Kingma & Ba, 2014) optimizer and polynomial learning rate scheduler to train ExLM, and we set the learning rate as 5e-4 and warmup steps as $10$K. The total training steps are $125$K and each batch has $2048$ samples at maximum. We also implement the ExLM model using the Fairseq library. For more pre-training hyper-parameters, please refer to Table 10.
@@
Figure 10: We visualize the model’s predictions when the input is “This is [MASK], and I’m very [MASK] to see this.” ($k=4$). The yellow nodes represent the expanded states corresponding to the first [MASK] token, while the brown nodes represent the expanded states corresponding to the second [MASK] token. The x-axis and y-axis show the top-3 predicted word for each graph node.
@@
  • ExLM successfully represents the different choices of each [MASK] token using distinct expanded states. Since $k=4$, the model predicts four different choices for both [MASK] tokens in the input. These four choices typically represent different aspects of the semantics. For example, the first [MASK] token’s choices include three categories: a noun (e.g., it, news), negative emotion words (e.g., terrible, awful), and positive emotion words (e.g., amazing, beautiful, wonderful), each represented by a different expanded state. This effectively avoids semantic ambiguity in the context, i.e., the semantic multimodality. Similarly, the second [MASK] token’s choices include positive emotions (e.g., glad, happy) and negative emotions (e.g., sorry), and these different types of semantic information are also represented by distinct expanded states.
@@
  • ExLM typically has lower uncertainty for [MASK] tokens that appear later. As shown in Figure 10, the word probability distributions of the four expanded states corresponding to the first [MASK] token are often more evenly distributed, meaning that the differences of word probabilities across different words in each state are not very large. However, for the four expanded states corresponding to the second [MASK] token, the model usually predicts one word with a probability significantly higher than the others. This phenomenon occurs due to the directionality of the DAG. Specifically, in ExLM’s DAG, transitions only occur from earlier states to later states, and the choices of earlier states have a greater impact on the final result, which means that the earlier states tend to have higher uncertainty.
@@
Q.1 Entropy Analysis on ExLM
We conduct an entropy analysis to compare the severity of the semantic multimodality under different mask ratios $p$ in the input context for ExLM and MLM. The results are shown in Figure 12. The results reveal that ExLM has significantly lower entropy in its prediction distribution than MLM, indicating lower uncertainty in dealing with ambiguous contexts. This further proves that ExLM significantly mitigates the negative impact of semantic multimodality and can still produce relatively confident predictions even when contextual semantic corruption is severe. Additionally, as $k$ increases, the uncertainty of ExLM decreases, suggesting that increasing $k$ expands the model’s semantic space, helping to alleviate the negative effects of the semantic multimodality. Moreover, under higher mask ratios, ExLM shows a more noticeable reduction in uncertainty compared to MLM, explaining why ExLM still maintains relatively strong performance under high mask ratios.
@@
We compare the training efficiency differences between MLM and ExLM ($k=4$) with the same pre-training configuration. Both models are trained on two Tesla A100 80G GPUs under the hyperparameters from Table 5, and their training cost statistics are shown in Table 12. From this table, we can know that the training cost of ExLM ($k=4$) is approximately 1.9 times that of the MLM with the same pre-training configuration. To compare the performance differences between MLM and ExLM ($k=4$) under equal training costs, we increase the training steps of MLM from $50{,}000$ to $100{,}000$, and refer to this model as Vanilla MLM++. The performance of Vanilla MLM++ is shown in Table 2, where we can see that although increasing the training steps significantly improves MLM performance, it still remains lower than that of ExLM. This further demonstrates that ExLM achieves stronger performance under the same training cost.